# An algebraic theory of Łojasiewicz exponents

### v1

@@ 5.6. Finite maximizers in the ideal case
Finally, the valuations coming from the codimension–one components of the exceptional locus of the normalized blowup $\operatorname{Proj}\overline{\mathcal{R}(\mathfrak{a})}\to\operatorname{Spec}R$ are precisely the classical Rees valuations of $\mathfrak{a}$; see [30, Chapter 10]. Therefore, the maximum above may be taken over the Rees valuations of $\mathfrak{a}$, as claimed. ∎
@@ 6.1. Hilbert–Samuel multiplicity and mixed multiplicities
$e_{i}({\mathfrak{a}_{\bullet}}^{[i]},{\mathfrak{b}_{\bullet}}^{[d-i]})=\limsup_{p\rightarrow\infty}\frac{e(\mathfrak{a}_{p}^{[i]},\mathfrak{b}_{p}^{[d-i]})}{n^{d}}.$
@@ 6.4. Mixed Łojasiewicz ratios and Teissier-type monotonicity
It follows that $\Lambda_{i}(\mathfrak{a}_{pt})\geq\Lambda_{i+1}(\mathfrak{a}_{pt})$. Dividing both sides of this inequality by appropriate powers of $pt$ and taking the limit along $p\to\infty$ then give $\Lambda_{1}({\mathfrak{a}_{\bullet}})\geq\Lambda_{2}({\mathfrak{a}_{\bullet}})\geq\cdots\geq\Lambda_{d}({\mathfrak{a}_{\bullet}})$ as claimed. ∎
@@ 7.1. Valuative definition of the log canonical threshold
where $v$ ranges over all divisorial valuations of $R$ centered at $\mathfrak{m}$. It was observed in [31] that, in general, there may not exist a divisorial valuation $v$ realizing $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$. On the other hand, it was also shown in [31, Theorem A], that given any graded family $\mathfrak{a}_{\bullet}$}, there is a real valuation $v$ of $R$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, i.e., $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}$.
@@
When ${\mathfrak{b}_{\bullet}}=\{\mathfrak{m}^{q}\}_{q\geq 1}$, we have $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq\operatorname{\mathrm{lct}}(\mathfrak{m})$. On the other hand, since $R$ is regular of dimension $d$, one has $\operatorname{\mathrm{lct}}(\mathfrak{m})=d$; see, for instance [35, § 9]. Thus, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq d$. ∎
@@ 7.3. Generic hyperplane sections and the Łojasiewicz exponent
    ([28, Theorem 2.1]) Integral closure is compatible with specialization by a generic element of $I$.
@@
    ([26, Proposition 3.7]) For a general element $x\in I$, one has
@@
Assume that $R$ satisfies the hypotheses of Remark 7.3. Let $\mathfrak{a}_{\bullet}$} be an $\mathfrak{m}$–primary filtration that is eventually degree-one normal, and set $I=\mathfrak{a}_{1}$. Then, there exists a Zariski open dense subset $U\subseteq I/\mathfrak{m}I$ such that for every $x\in I$ with class in $U$, and every integer $p\geq 1$, there exists an integer $t_{0}=t_{0}(p)$ with the property that
@@
Let $x\in I$ be general as in Remark 7.3(ii) and set $R^{\prime}=R/(x)$. By [26, Proposition 3.7], there exists $n_{0}$ such that
@@ 8.1. Rigidity I: the common extremal valuation principle
Only if. Assume $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})$. Choose a real valuation $v_{0}$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, whose existence was established in [31, Theorem A]. That is, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=A(v_{0})/v_{0}({\mathfrak{a}_{\bullet}})$. By the valuative definition of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ in Theorem 5.6, we have
@@ 8.2. Rigidity II: the monomial case
Since $I$ is monomial, the Rees valuations of $I$ are precisely the monomial valuations obtained from the (non-redundant) bounding hyperplanes of the Newton polyhedron $\operatorname{NP}(I)$; see [30, Theorem 10.3.5]. For an $\mathfrak{m}$–primary monomial ideal, the relevant bounding hyperplanes are exactly those defining the compact facets of $\operatorname{NP}(I)$, hence they are indexed by the primitive inward normal vectors $u\in\mathcal{U}(I)$, and the corresponding valuations are the $v_{u}$. Therefore
@@
If $\Theta(I)=1$, then $\operatorname{\mathrm{lct}}(I)\cdot\mathcal{L}_{\mathfrak{m}}(I)=d=\operatorname{\mathrm{lct}}(\mathfrak{m})$. By Theorem 8.1, equality forces a common valuation computing both $\operatorname{\mathrm{lct}}(I)$ and $\mathcal{L}_{\mathfrak{m}}(I)$ and also satisfying $A(v)/v(\mathfrak{m})=d$. In the monomial category, such a valuation must be $v_{u}$ for some $u\in\mathbb{Z}_{>0}^{d}$, and then
@@ 9.3. Stratification by the computing valuation
These two ratios are equal precisely at $t=\frac{1}{3}$. Consequently,
@@
Let $T$ be a Noetherian scheme, let $X$ be a Noetherian integral scheme, and set $X_{T}:=X\times T$. Let $\pi:Y\to X_{T}$ be a proper birational morphism with $Y$ normal, let $E\subseteq Y$ be a prime divisor whose image dominates $T$, and let $\mathcal{I}\subseteq\mathcal{O}_{X_{T}}$ be a coherent ideal sheaf, and for $t\in T$ let $\mathcal{I}_{t}$ be its fiber on $X\times\{t\}$. Define
@@ 9.4. Geodesics of filtrations and fractional linearity of $1/L$
In particular, $t\mapsto 1/L(t)$ is fractional linear (hence $C^{\infty}$) on each such $I$, and its second derivative has constant sign on $I$. Moreover, on a given component $I$ the restriction $t\mapsto 1/L(t)$ is affine if and only if the affine functions $\alpha_{j}$ and $\beta_{j}$ are proportional on $I$.
@@
whose sign is constant on $I$ because the denominator is positive and the numerator is constant. Finally, $\beta_{j}/\alpha_{j}$ is affine on $I$ if and only if $b_{1}a_{0}-b_{0}a_{1}=0$, i.e. if and only if $\alpha_{j}$ and $\beta_{j}$ are proportional affine functions on $I$. ∎
@@ 10.4. Integral closure and Newton polyhedra in toric rings
Finally, $\operatorname{NP}(I)$ is a rational polyhedron, so it is the intersection of its supporting halfspaces defined by the primitive inward normals of its compact faces. Therefore, the above inequalities are equivalent to $m\in\operatorname{NP}(I)\cap S$. ∎
@@ 10.6. Facet normals and finite candidate sets
Note that, in Theorem 10.11, only compact facets contribute since $\mathfrak{a}$ being $\mathfrak{m}_{S}$–primary implies $h_{\operatorname{NP}(\mathfrak{a})}(u)>0$ precisely along inward normals to compact faces.
@@ 10.7. Monomial filtrations in the toric setting
Let $\sigma\subseteq N_{\mathbb{R}}$ be a strongly convex rational polyhedral cone and set $\sigma^{\circ}$ for its interior. Let $\Sigma$ be a finite rational polyhedral fan supported on $\sigma$ with rays $\rho_{1},\dots,\rho_{M}$ and primitive generators $u_{1},\dots,u_{M}\in N$. Let $\alpha,\beta:\sigma^{\circ}\to\mathbb{R}_{>0}$ be functions such that:
@@
    $\alpha$ and $\beta$ extend to continuous functions on $\sigma^{\circ}$ and are positive on $\sigma^{\circ}$;
@@
    for every cone $\tau\in\Sigma$, the restrictions $\alpha|_{\tau^{\circ}}$ and $\beta|_{\tau^{\circ}}$ are linear forms in $u$.
Then,
$\sup_{u\in\sigma^{\circ}\cap N}\frac{\alpha(u)}{\beta(u)}\;=\;\max_{1\leq j\leq M}\frac{\alpha(u_{j})}{\beta(u_{j})}.$
In particular, the supremum is a maximum attained at a primitive ray generator of $\Sigma$.
@@
Since $\beta(u)>0$ on $\sigma^{\circ}$, we may normalize on the compact slice
$K:=\{u\in\sigma^{\circ}:\beta(u)=1\}.$
On each cone $\tau\in\Sigma$, the set $K\cap\tau$ is a compact polytope and $\alpha/\beta=\alpha$ on $K$ is linear on $\tau$ by hypothesis. Hence, the maximum of $\alpha$ on $K\cap\tau$ is attained at a vertex of that polytope, which lies on an extreme ray of $\tau$. Taking the maximum over finitely many cones yields that the global supremum on $K$ is attained at one of the finitely many ray generators $u_{1},\dots,u_{M}$ of $\Sigma$, giving the claimed formula. ∎
Corollary 10.16 (Polyhedral Newton data $\Rightarrow$ finite toric candidate set).
Assume that the toric valuation functions
$u\longmapsto v_{u}(\mathfrak{a}_{\bullet}),\quad u\longmapsto v_{u}(\mathfrak{b}_{\bullet})$
are piecewise-linear on a common finite fan $\Sigma$ supported on $\sigma$ (for example, when the relevant Newton data is rational polyhedral and one takes $\Sigma$ a common refinement of the associated normal fans). Then, the toric supremum is a finite maximum:
$\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}\;=\;\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$
@@
Apply Lemma 10.15 with $\alpha(u)=v_{u}(a_{\bullet})$ and $\beta(u)=v_{u}(b_{\bullet})$. ∎
@@
Let $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and $\mathfrak{b}_{\bullet}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be graded families of monomial ideals in $R$.
(1) (Universal toric lower bound) One has
$\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})\ \geq\ \sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}=\sup_{u\in\sigma^{\circ}\cap N}\frac{\inf_{y\in\operatorname{NP}(\mathfrak{a}_{\bullet})}\langle u,y\rangle}{\inf_{x\in\operatorname{NP}(\mathfrak{b}_{\bullet})}\langle u,x\rangle}.$
(2) (Finite–max among toric valuations) Suppose that the functions $u\mapsto v_{u}(\mathfrak{a}_{\bullet})$ and $u\mapsto v_{u}(\mathfrak{b}_{\bullet})$ are piecewise-linear on a common finite rational fan $\Sigma$ supported on $\sigma$. Then
$\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}=\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$
(3) (Equality with $\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})$ under toric testing) Assume in addition that for all $p,q$ the containment $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$ is equivalent to the system of inequalities
$v_{u}(\mathfrak{b}_{q})\geq v_{u}(\mathfrak{a}_{p})\quad\text{for all }u\in\mathrm{Ray}(\Sigma).$
Then,
$\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})=\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$

### v2

@@ 5.6. Finite maximizers in the ideal case
Finally, the valuations coming from the codimension–one components of the exceptional locus of the normalized blowup $\operatorname{Proj}\overline{\mathcal{R}(\mathfrak{a})}\to\operatorname{Spec}R$ are precisely the classical Rees valuations of $\mathfrak{a}$; see [34, Chapter 10]. Therefore, the maximum above may be taken over the Rees valuations of $\mathfrak{a}$, as claimed. ∎
@@ 6.1. Hilbert–Samuel multiplicity and mixed multiplicities
$e_{i}({\mathfrak{a}_{\bullet}}^{[i]},{\mathfrak{b}_{\bullet}}^{[d-i]})=\limsup_{p\rightarrow\infty}\frac{e(\mathfrak{a}_{p}^{[i]},\mathfrak{b}_{p}^{[d-i]})}{p^{d}}.$
@@ 6.4. Mixed Łojasiewicz ratios and Teissier-type monotonicity
It follows that Dividing both sides of this inequality by appropriate powers of $pt$ and taking the limit along $p\to\infty$ then give $\Lambda_{1}({\mathfrak{a}_{\bullet}})\geq\Lambda_{2}({\mathfrak{a}_{\bullet}})\geq\cdots\geq\Lambda_{d}({\mathfrak{a}_{\bullet}})$ as claimed. ∎
@@ 7.1. Valuative definition of the log canonical threshold
where $v$ ranges over all divisorial valuations of $R$ centered at $\mathfrak{m}$. It was observed in [35] that, in general, there may not exist a divisorial valuation $v$ realizing $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$. On the other hand, it was also shown in [35, Theorem A], that given any graded family $\mathfrak{a}_{\bullet}$}, there is a real valuation $v$ of $R$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, i.e., $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}$.
@@
When ${\mathfrak{b}_{\bullet}}=\{\mathfrak{m}^{q}\}_{q\geq 1}$, we have $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq\operatorname{\mathrm{lct}}(\mathfrak{m})$. On the other hand, since $R$ is regular of dimension $d$, one has $\operatorname{\mathrm{lct}}(\mathfrak{m})=d$; see, for instance [39, § 9]. Thus, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq d$. ∎
@@ 7.3. Generic hyperplane sections and the Łojasiewicz exponent
    ([32, Theorem 2.1]) Integral closure is compatible with specialization by a generic element of $I$.
@@
    ([30, Proposition 3.7]) For a general element $x\in I$, one has
@@
Assume that $R$ satisfies the hypotheses of Remark 7.3 and $\dim R\geq 2$. Let $\mathfrak{a}_{\bullet}$} be an $\mathfrak{m}$–primary filtration that is eventually degree-one normal, and set $I=\mathfrak{a}_{1}$. Then, there exists a Zariski open dense subset $U\subseteq I/\mathfrak{m}I$ such that for every $x\in I$ with class in $U$, and every integer $p\geq 1$, there exists an integer $t_{0}=t_{0}(p)$ with the property that
@@
Let $x\in I$ be general as in Remark 7.3(ii) and set $R^{\prime}=R/(x)$. By [30, Proposition 3.7], there exists $n_{0}$ such that
@@ 8.1. Rigidity I: the common extremal valuation principle
Only if. Assume $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})$. Choose a real valuation $v_{0}$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, whose existence was established in [35, Theorem A]. That is, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=A(v_{0})/v_{0}({\mathfrak{a}_{\bullet}})$. By the valuative definition of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ in Theorem 5.6, we have
@@ 8.2. Rigidity II: the monomial case
Since $I$ is monomial, the Rees valuations of $I$ are precisely the monomial valuations obtained from the (non-redundant) bounding hyperplanes of the Newton polyhedron $\operatorname{NP}(I)$; see [34, Theorem 10.3.5]. For an $\mathfrak{m}$–primary monomial ideal, the relevant bounding hyperplanes are exactly those defining the compact facets of $\operatorname{NP}(I)$, hence they are indexed by the primitive inward normal vectors $u\in\mathcal{U}(I)$, and the corresponding valuations are the $v_{u}$. Therefore
@@
Consider the equality $\Theta(I)=1$, i.e., $\operatorname{\mathrm{lct}}(I)\cdot\mathcal{L}_{\mathfrak{m}}(I)=d=\operatorname{\mathrm{lct}}(\mathfrak{m})$. Let $v_{u}$, for some $u\in\mathbb{Z}^{d}_{>0}$, be a monomial valuation that computes $\operatorname{\mathrm{lct}}(I)$ as in Step 1. Then, since $\mathcal{L}_{\mathfrak{m}}(I)\geq\frac{v_{u}(I)}{v_{u}(\mathfrak{m})}$ and $\frac{A(v_{u})}{v_{u}(\mathfrak{m})}\geq\operatorname{\mathrm{lct}}(\mathfrak{m})$, the equality $\Theta(I)=1$ forces $v_{u}$ to compute both $\mathcal{L}_{\mathfrak{m}}(I)$ and $\operatorname{\mathrm{lct}}(\mathfrak{m})$. Thus,
@@ 9.3. Stratification by the computing valuation
These two ratios are equal precisely at $t=1\frac{1}{3}$. Consequently,
@@
Let $T$ be a Noetherian scheme, let $X$ be a Noetherian integral scheme, and set $X_{T}:=X\times T$. Let $\pi:Y\to X_{T}$ be a proper birational morphism with $Y$ normal and $Y\to T$ a flat family, and let $E\subseteq Y$ be an effective Cartier divisor that dominates $T$. Let $\mathcal{I}\subseteq\mathcal{O}_{X_{T}}$ be a coherent ideal sheaf, and for $t\in T$ let $\mathcal{I}_{t}$ be its fiber on $X\times\{t\}$. Define
@@ 9.4. Geodesics of filtrations and fractional linearity of $1/L$
In particular, $t\mapsto 1/L(t)$ is fractional linear (hence $C^{\infty}$) on each such $I$, and its second derivative has constant sign on $I$. Moreover, on a given component $I$ the restriction $t\mapsto 1/L(t)$ is affine if and only if either $\alpha_{j}$ is constant on $I$ or the affine functions $\alpha_{j}$ and $\beta_{j}$ are proportional on $I$.
@@
whose sign is constant on $I$ because the denominator is positive and the numerator is constant. Finally, $\beta_{j}/\alpha_{j}$ is affine on $I$ if and only if $a_{1}(b_{1}a_{0}-b_{0}a_{1})=0$, i.e. if and only if either $\alpha_{j}$ is constant or $\alpha_{j}$ and $\beta_{j}$ are proportional affine functions on $I$. ∎
@@ 10.4. Integral closure and Newton polyhedra in toric rings
Finally, $\operatorname{NP}(I)$ is a rational polyhedron, so it is the intersection of its supporting halfspaces defined by the primitive inward normals of its compact facets. Therefore, the above inequalities are equivalent to $m\in\operatorname{NP}(I)\cap S$. ∎
@@ 10.7. Monomial filtrations in the toric setting
Let $\alpha,\beta:\sigma\to\mathbb{R}_{\geq 0}$ be continuous functions that are positively homogeneous of degree $1$. Assume that:
@@
    $\beta(u)>0$ for all $u\in\sigma^{\circ}$;
@@
    there exists a finite rational polyhedral fan $\Sigma$ supported on $\sigma$ such that $\alpha$ and $\beta$ are linear on every cone $\tau\in\Sigma$.
Let $\rho_{1},\dots,\rho_{M}$ be the rays of $\Sigma$ and $u_{1},\dots,u_{M}\in N$ their primitive generators. Then,
$\sup_{u\in\sigma^{\circ}\cap N}\frac{\alpha(u)}{\beta(u)}=\sup_{u\in\sigma^{\circ}}\frac{\alpha(u)}{\beta(u)}=\max_{\{j:\beta(u_{j})>0\}}\frac{\alpha(u_{j})}{\beta(u_{j})}.$
@@
Fix a cone $\tau\in\Sigma$, and let $w_{1},\dots,w_{s}$ be the primitive generators of its rays. Every $u\in\tau$ can be written as $u=\sum_{i}\lambda_{i}w_{i}$ with $\lambda_{i}\geq 0$. Since $\alpha$ and $\beta$ are linear on $\tau$, we have $\alpha(u)=\sum_{i}\lambda_{i}\alpha(w_{i})$ and $\beta(u)=\sum_{i}\lambda_{i}\beta(w_{i})$.
If $u\in\tau^{\circ}$, then $\beta(u)>0$ by assumption. Hence,
$\frac{\alpha(u)}{\beta(u)}=\frac{\sum_{i}\lambda_{i}\alpha(w_{i})}{\sum_{i}\lambda_{i}\beta(w_{i})}=\sum_{\beta(w_{i})>0}\frac{\lambda_{i}\beta(w_{i})}{\sum_{k}\lambda_{k}\beta(w_{k})}\cdot\frac{\alpha(w_{i})}{\beta(w_{i})}.$
Thus $\alpha(u)/\beta(u)$ is a convex combination of the finitely many numbers $\alpha(w_{i})/\beta(w_{i})$ with $\beta(w_{i})>0$. Therefore,
$\frac{\alpha(u)}{\beta(u)}\leq\max_{\{i:\beta(w_{i})>0\}}\frac{\alpha(w_{i})}{\beta(w_{i})}$
for every $u\in\tau^{\circ}$, and hence the same bound holds for $\sup_{u\in\tau^{\circ}}$.
Conversely, let $w_{i}$ be a ray generator of $\tau$ with $\beta(w_{i})>0$. Choose $z\in\tau^{\circ}$ and set $u_{\varepsilon}=(1-\varepsilon)w_{i}+\varepsilon z$ for $0<\varepsilon<1$. Then, $u_{\varepsilon}\in\tau^{\circ}$ and $u_{\varepsilon}\to w_{i}$ as $\varepsilon\to 0^{+}$. By continuity of $\alpha$ and $\beta$ and since $\beta(w_{i})>0$, we have $\alpha(u_{\varepsilon})/\beta(u_{\varepsilon})\to\alpha(w_{i})/\beta(w_{i})$. Hence, $\sup_{u\in\tau^{\circ}}\alpha(u)/\beta(u)\geq\alpha(w_{i})/\beta(w_{i})$.
Combining the two inequalities shows that
$\sup_{u\in\tau^{\circ}}\frac{\alpha(u)}{\beta(u)}=\max_{\{i:\beta(w_{i})>0\}}\frac{\alpha(w_{i})}{\beta(w_{i})}.$
Since $\Sigma$ has finitely many cones and every ray of $\Sigma$ occurs among the rays of some cone, taking the maximum over all cones gives
$\sup_{u\in\sigma^{\circ}}\frac{\alpha(u)}{\beta(u)}=\max_{\{j:\beta(u_{j})>0\}}\frac{\alpha(u_{j})}{\beta(u_{j})}.$
Finally, because $\alpha/\beta$ is homogeneous of degree 0, it suffices to approximate interior points by interior lattice points. Rational points are dense in $\sigma^{\circ}$, and any interior rational point becomes a lattice point after multiplying by a positive integer. Homogeneity implies that the ratio $\alpha/\beta$ is unchanged by such scaling, so the supremum over $\sigma^{\circ}$ equals the supremum over $\sigma^{\circ}\cap N$. ∎
In the next few results, if for each monomial filtration $\mathfrak{a}_{\bullet}$}, the function $u\mapsto v_{u}({\mathfrak{a}_{\bullet}})$ on $\sigma^{\circ}\cap N$ admits a continuous positively homogeneous extension to $\sigma$, then we shall denote the value of this extension at $u\in\sigma$ by $\widetilde{v_{u}}({\mathfrak{a}_{\bullet}})$.
Corollary 10.16.
Let $\mathfrak{a}_{\bullet}$} and $\mathfrak{b}_{\bullet}$} be monomial filtrations. Assume that the functions $\sigma^{\circ}\cap N\to\mathbb{R}_{\geq 0}$, $u\mapsto v_{u}({\mathfrak{a}_{\bullet}})$ and $u\mapsto v_{u}({\mathfrak{b}_{\bullet}})$, admit continuous extensions to $\sigma$, positively homogeneous of degree $1$, and that there exists a finite rational polyhedral fan $\Sigma$ supported on $\sigma$ such that both extended functions are linear on every cone $\tau\in\Sigma$. Assume also that $v_{u}({\mathfrak{b}_{\bullet}})>0$ for all $u\in\sigma^{\circ}$.
Let $\rho_{1},\dots,\rho_{M}$ be the rays of $\Sigma$ and let $u_{1},\dots,u_{M}\in N$ be their primitive generators. Then,
$\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}=\max_{\{j:\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})>0\}}\frac{\widetilde{v}_{u_{j}}({\mathfrak{a}_{\bullet}})}{\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})}.$
@@
Apply Lemma 10.15 with $\alpha(u)=\widetilde{v}_{u}({\mathfrak{a}_{\bullet}})$ and $\beta(u)=\widetilde{v}_{u}({\mathfrak{b}_{\bullet}})$. Since $\widetilde{v}_{u}({\mathfrak{a}_{\bullet}})=v_{u}({\mathfrak{a}_{\bullet}})$ and $\widetilde{v}_{u}({\mathfrak{b}_{\bullet}})=v_{u}({\mathfrak{b}_{\bullet}})$ for $u\in\sigma^{\circ}\cap N$, the claimed equality follows. ∎
@@
Let $\mathfrak{a}_{\bullet}$} and $\mathfrak{b}_{\bullet}$} be $\mathfrak{m}_{S}$-primary monomial filtrations, and assume that $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is Noetherian. Suppose moreover that there exist an $\mathfrak{m}_{S}$-primary monomial ideal $\mathfrak{q}$ and a finite rational polyhedral fan $\Sigma$ supported on $\sigma$ such that:
  1. (1)
    $\mathfrak{a}_{\bullet}$} and $\mathfrak{b}_{\bullet}$} are linearly bounded with respect to $\mathfrak{q}$;
  2. (2)
    the functions $u\mapsto v_{u}({\mathfrak{a}_{\bullet}})$ and $u\mapsto v_{u}({\mathfrak{b}_{\bullet}})$ on $\sigma^{\circ}\cap N\to\mathbb{R}_{\geq 0}$ admit continuous extensions to $\sigma$, positively homogeneous of degree $1$, and linear on every cone $\tau\in\Sigma$.
Let $\rho_{1},\dots,\rho_{M}$ be the rays of $\Sigma$, and let $u_{1},\dots,u_{M}\in N$ be their primitive generators. Then
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}=\max_{\{j:\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})>0\}}\frac{\widetilde{v}_{u_{j}}({\mathfrak{a}_{\bullet}})}{\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})}.$
In particular, $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is computed by finitely many toric divisorial valuations.